OpenStudy (anonymous):

(4p^(2/3))/(4p)^2 - solve for p

zepdrix (zepdrix):

Hey Meg :) We can not "solve" as there is no equality sign present. I assume you mean simplify.\[\large\rm \frac{4p^{2/3}}{(4p)^2}\]

zepdrix (zepdrix):

Let's start in the denominator. Recall that if you're squaring a quantity in brackets, you have to square each and every item inside, \(\large\rm (ab)^2=a^2b^2\)

zepdrix (zepdrix):

So in the denominator you're getting \(\large\rm (4p)^2=4^2p^2=16p^2\) ya?